# Decom-Renorm-Merge: Model Merging on the Right Space Improves Multitasking

### v1

@@ Abstract
In the era of large-scale training, model merging has evolved into a tool for creating multitasking models efficiently. It enables the knowledge of models to be fused, without the need for heavy computation as required in traditional multitask learning. Existing merging methods often assume that entries at identical positions in weight matrices serve the same function, enabling straightforward entry-wise comparison and merging. However, this assumption overlooks the complexity of finetuned neural networks, where neurons may develop distinct feature compositions, making direct entry-wise merging problematic. We present Decom-Renorm-Merge (DRM), a simple yet effective approach that leverages Singular Value Decomposition to decompose and coordinate weight matrices into an aligned joint space, where entry-wise merging becomes possible. We showcase the effectiveness of DRM across various settings ranging from smaller encoder-based such as ViT and DeBERTa, encoder-decoder-based such as T5, and larger decoder-based such as Llama3.1-8B. Our experimental results show that DRM outperforms several state-of-the-art merging techniques across full finetuning and low-rank adaptation settings. Moreover, our analysis reveals renormalization as the crucial component for creating a robust and even joint space for merging, significantly contributing to the method’s performance.
@@ 1 Introduction
Parameter interference is identified in TIES Merging [83] as a key obstacle to effective model merging. They demonstrate that two major interferences are redundant parameter values and sign disagreements at corresponding positions in the weight matrices. Yadav et al. [83] propose to address the issue by first pruning small entries from the task vectors. Then, in each position, one dominant sign is selected, and all other entries with opposite sign are zeroed out. The pruned entries are ignored during the final averaging step.
These interference mitigation operations, employed in TIES [83] and its variants [85], rely on a crucial assumption: task vector entries at identical positions serve consistent semantic roles and functions. This assumption allows entry-wise comparison, pruning, dominant sign selection, and the subsequent averaging to be done. However, this fundamental assumption may not hold in practice, as: (a) finetuning can lead to changes in feature arrangement of the activations and (b) a single neuron often carries not just a single, but a composition of several overlapping features, a phenomenon known as polysemanticity [19, 67, 53]. Based on these observations, we formulate our central hypothesis: for entry-wise interference reduction strategies to be valid during model merging, task vectors must first be decomposed and then coordinated into a shared representation space.
@@
Figure 1: Decom-Renorm-Merge (DRM) is a model-merging method for building multitask models. Different models may not share the same weight parameterization. Thus, merging should occur in a shared decomposed weight space, not the original parameter space. DRM merges models’ weight deltas $\Delta W^{(t)}$—the difference of each finetuned model from a shared base model—into a single merged delta. DRM consists of four main steps: (a) Decompose: Concatenate the $\Delta W^{(t)}$ matrices horizontally (or vertically), then apply SVD to decompose them into a shared basis $U$ and individual weights $V_{t}$. Although the combined $V$ is orthonormal, each individual split $V_{t}$ is not. (b) Renormalize each row vector $v_{t,i}$ of $V_{t}$ to unit length, and scale the corresponding singular value to preserve the magnitude. This can be viewed as compensating for $V_{t}$’s non-orthonormality as illustrated in (c). (d) Prune each renormalized individual singular vector matrix $\tilde{V}_{t}$ by keeping only the entries within top-$k$magnitudes. (e) Merge the pruned singular vector matrices across models using sign election and disjoint averaging.
We demonstrate the effectiveness of DRM across various setups, including: (1) Diverse model architectures and sizes. Ranging from smaller encoder-based ViT-B/32, ViT-L/14 [15, 61], and DeBERTa-Base [26]; mid-sized, encoder-decoder model T5-Base and T5-Large [63]; to large decoder-based Llama3.1-8B [24]. (2) Comprehensive task categories and modalities. Including image recognition and natural language understanding. (3) With and without validation set for merging techniques’ hyperparameter tuning. Notably, even with no validation set available, DRM outperforms the strongest baselines by $5.0\%$ and $1.9\%$ in merging ViT-B/32 and ViT-L/14, respectively, $9.3\%$ on DeBERTa-Base-v3, and $1.9\%$ on low-rank (LoRA) adapted Llama3.1-8B, on absolute scales. Additionally, we demonstrate that DRM has better performance retention as the number of tasks being merged increases, compared to other baselines.
@@
  • We identify and analyze the critical role of renormalization (Section 6) in enabling stable and robust merging within the shared decomposed weight space.
@@ 3 Motivation: Merging models on the right space
3.1 Different finetuned models may not share the same weight parameterizations
@@
Although, under the infinite width assumption, neural networks follow a deterministic trajectory in functional space during training [37], this does not guarantee a unique solution in parameter space. In fact, different model parameterizations may lead to functionally equivalent models. These weight symmetries are supported by empirical findings in weight permutation studies [58, 54, 74, 20, 3, 39, 72], which suggests that even under the same pretrained feature space, functionally similar features can be manipulated into different arrangements throughout the course of finetuning.
@@ 4.1 Hypothesis
Building on the above observations, we know that weight delta entries at identical positions do not necessarily encode consistent semantics or functionalities. This invalidates direct entry-wise interference reduction of the weight delta matrices. We make the following hypothesis: weight deltas of differently adapted models must be decomposed and then coordinated into an identical representation space, before proper interference reduced merging can be carried out.
@@ 4.2 Algorithm
With the hypothesis in mind, we propose Decom-Renorm-Merge (DRM), comprises (1) shared decomposition via SVD, for aligning different weight deltas into a joint representation space, (2) per-task renormalization for stabilizing each individual basis and enabling downstream interference reductions, (3) TIES’ operations applied in the Decompose-renormalized joint space, including pruning, sign election, and disjoint averaging. See Appendix E for analysis on different decompositions.
@@
4) Pruning: For each task $t$, we keep only the top-$k\%$ largest-magnitude entries in $\tilde{V}^{T}_{t}$, and zero out the others. It is vital that this step be done after renormalization to ensure stability, pruning on unnormalized magnitudes could distort the basis vectors. We set the default to top-$20\%$
5) Sign Election: To resolve inter-task sign conflicts, we compute a sign mask, $S=\operatorname{sgn}(\sum_{t=1}^{N}\Sigma_{t}\tilde{V}_{t}^{T})$, selecting a magnitude dominant sign for each position. Then, we zero out entries with signs opposite to the elected mask. This procedure is done in scale-variant fashion by multiplying the $\Sigma_{t}$ in.
@@
Here, $\lambda_{t}$ is a task-specific scalar weighting coefficient, with default $\lambda=1.0$ for all tasks. $\Gamma$ is an entry-wise averaging reciprocal mask that ensures averaging procedure is performed only over non-zero entries. We define $\Gamma\in\mathbb{R}^{m\times n}$ with entry at row $i$, column $j$ as the following reciprocal:
@@ Methods
For merging of large language models, we adapt Llama3.1-8B [76] to five GLUE tasks individually: MNLI [80], QNLI [78], RTE [78], COLA [79], and SST2 [70] For this setup, we perform low-rank adaptation ($r=16$) instead of full finetuning as done previously. Table 2 shows that without validation data (no hyperparameter tuning), DRM-H outperforms other competitors by more than $1.9\%$. In the presence of hyperparameter tuning, DRM-V outperforms the strongest baseline by a margin of $3.9\%$. Notice that, unlike in the previous winning scenarios, where DRM-V is usually subpar or relatively close to DRM-H. Reversed here, DRM-V with hyperparameter tuning surpasses DRM-H by a margin of $1.0\%$. We provide analysis on this phenomenon in Appendix G.
@@ 6 Understanding the Decompose-Renormalized Joint Space
Seeing the empirical verification of the DRM, we hope to understand more about our method, particularly the intrinsic nature of the Decompose-renormalized joint space used for the merging.
We perform a series of studies to answer the following questions: 1) Why is it crucial that magnitude pruning be done on renormalized singular vector matrices? Shouldn’t pruning on unrenormalized basis vector matrices be even better? Since they better capture the real scale of each basis entry. 2) How does reducing interference on the Decompose-renormalized space reflect interference on the original parameter space? Does reducing interference in one space directly influence the other?
@@
6.1 Why is it Crucial that Weight Pruning be Done on Renormalized Singular Vector Matrices?
After partitioning the right singular vector matrix $V^{T}$ from the concatenated SVD, the row basis vectors lose the orthonormal property. We renormalize each row to unit norm [^1] Unlike the unit length property, orthogonality of the split matrices, is not easily reattainable. We observe that enforcing orthogonality through Gram-Schmidt process harshly degrades the performance. before applying interference reduction techniques. As shown in Section 5.3, renormalization is central to DRM. We present the investigation of effect of renormalization on the pruning behavior in Figure 3. It displays the percentages of entries being dropped from each basis vector. We see that without renormalization, the dropping densities oscillate severely, with some vectors have almost all of their entries zeroed out. Removing top principle components like this would quickly deteriorate downstream performance. By adding in renormalization, the dropping densities are stabilized evenly across all vectors, bounded to around the configured prune rate of 50%, no single vector will be wholly pruned.
Table 4: Average sign agreement after pruning different spaces. We compare between pruning in the original parameter space (labeled Original Weight), versus pruning in the joint singular space. For the latter, sign agreement is measured in both the joint singular space (labeled DS), and after projecting back to the original space (labeled OS). Results demonstrate that reducing interference on the joint space does not correspond to higher sign agreement in the original parameter space. Note that the minimum sign agreement is $50\%$, when there are an equal number of positive and negative entries.
@@ 6.2 How Does Pruning on Decompose-renormalized Space Reflect Interference in Original Parameter Space?
However, our analysis in Table 4 reveals that although pruning in the Decompose-renormalized space effectively improves merging performance, it does not result in higher sign agreement when projected back to the original parameter space. Specifically, we first project the weight deltas into the Decompose-renormalized space, apply pruning to the basis entries. We then transform the results back to the original space and measure sign agreement, but we found no improvement.
Nevertheless, DRM still achieves mostly superior merging performance compared to methods reducing interference directly in the original parameter space. This suggests that interference reduction techniques lead to the most effective merging when applied in the Decompose-renormalized space, rather than directly on individual parameter entries. See Appendix F for more detailed results and comparison with no pruning applied.
@@ 8 Acknowledgements
We are grateful to Artit Suwanbandit and Patawee Prakrankamanant for inspiration and insightful discussions early into this work.
@@ Appendix C Significance Analysis of the Merging Performance
Appendix D Theoretical Bound of Difference Between Individual Task and the Concatenated Matrix
This section theoretically evaluates the bounded difference between weight delta of each individual task, and the horizontally concatenated matrix used in DRM-H. We express the difference through their corresponding singular values (i.e. $i$-th singular value of a single weight delta matrix, and $i$-th singular value of the concatenated matrix).
We begin by stating a theorem proven by Shamrai [68]:

### v2

@@ Abstract
In the era of large-scale training, model merging has evolved into a tool for creating multitasking models efficiently. It enables the knowledge of models to be fused, without the need for heavy computation as required in traditional multitask learning. Existing merging methods often assume that entries at identical positions in weight matrices serve the same function, enabling straightforward entry-wise comparison and merging. However, this assumption overlooks the complexity of finetuned neural networks, where neurons may develop distinct feature compositions, making direct entry-wise merging problematic. We present Decom-Renorm-Merge (DRM), a simple yet effective approach that leverages Singular Value Decomposition to decompose and coordinate weight matrices into an aligned joint space, where entry-wise merging becomes possible. We showcase the effectiveness of DRM across various settings ranging from smaller encoder-based such as ViT and DeBERTa, encoder-decoder-based such as T5, and larger decoder-based such as Llama3.1-8B. Our experimental results show that DRM outperforms several state-of-the-art merging techniques across full finetuning and low-rank adaptation settings. Moreover, our analysis reveals renormalization as the crucial component for creating a robust and even joint space for merging, significantly contributing to the method’s performance. [^1] Code and models are available at https://github.com/yophis/decom-renorm-merge.
@@ 1 Introduction
Parameter interference is identified in TIES Merging [83] as a key obstacle to an effective model merging. They demonstrate that two major interferences are redundant parameter values and sign disagreements at corresponding positions in the weight matrices. Yadav et al. [83] propose to address the issue by first pruning small entries from the task vectors. Then, in each position, one dominant sign is selected, and all other entries with opposite sign are zeroed out. The pruned entries are ignored during the final averaging step.
These interference mitigation operations, employed in TIES [83] and its variants [85], rely on a crucial assumption: task vector entries at identical positions serve consistent semantic roles and functions. This assumption allows entry-wise comparison, pruning, dominant sign selection, and the subsequent averaging to be done. However, this fundamental assumption may not hold in practice, as: (a) finetuning can lead to changes in feature arrangement of the activations [58, 20, 3], and (b) a single neuron often carries not just a single, but a composition of several overlapping features, a phenomenon known as polysemanticity [19, 67, 53]. Based on these observations, we formulate our central hypothesis: for entry-wise interference reduction strategies to be valid during model merging, task vectors must first be decomposed and then coordinated into a shared representation space.
@@
Figure 1: Decom-Renorm-Merge (DRM) is a model-merging method for building multitask models. Different models may not share the same weight parameterization. Thus, merging should occur in a shared decomposed weight space, not the original parameter space. DRM merges models’ weight deltas $\Delta W^{(t)}$—the difference of each finetuned model from a shared base model—into a single merged delta. DRM consists of four main steps: (a) Decompose: Concatenate the $\Delta W^{(t)}$ matrices horizontally (or vertically), then apply SVD to decompose them into a shared basis $U$ and individual weights $V_{t}$. Although the combined $V$ is orthonormal, each individual split $V_{t}$ is not. (b) Renormalize each row vector $v_{t,i}$ of $V_{t}$ to unit length, and scale the corresponding singular value to preserve the magnitude. This can be viewed as compensating for $V_{t}$’s non-orthonormality as illustrated in (c). (d) Prune each renormalized individual singular vector matrix $\tilde{V}_{t}$ by keeping only the entries within top-kk% magnitudes. (e) Merge the pruned singular vector matrices across models using sign election and disjoint averaging.
We demonstrate the effectiveness of DRM across various setups, including: (1) Diverse model architectures and sizes. Ranging from smaller encoder-based ViT-B/32, ViT-L/14 [15, 61], and DeBERTa-Base [26]; mid-sized, encoder-decoder model T5-Base and T5-Large [63]; to large decoder-based Llama3.1-8B [24]. (2) Comprehensive task categories and modalities, including image recognition and natural language understanding. (3) With and without validation set for merging techniques’ hyperparameter tuning. Notably, even with no validation set available, DRM outperforms the strongest baselines by $5.0\%$ and $1.9\%$ in merging ViT-B/32 and ViT-L/14, respectively, $9.3\%$ on DeBERTa-Base-v3, and $1.9\%$ on low-rank (LoRA) adapted Llama3.1-8B, on absolute scales. Additionally, we demonstrate that DRM has better performance retention as the number of tasks being merged increases, compared to other baselines.
@@
  • We identify the critical role of renormalization (Section 6) in enabling stable and robust merging within the shared decomposed weight space, through a series of empirical and theoretical analyses.
@@ 3 Motivation: Merging models on the right space
3.1 Different Finetuned Models may not Share the Same Weight Parameterization
@@
Although, under the infinite width assumption, neural networks follow a deterministic trajectory in functional space during training [37], this does not guarantee a unique solution in parameter space. In fact, different model parameterizations may lead to functionally equivalent models. These weight symmetries are supported by empirical findings in weight permutation studies [58, 54, 74, 20, 3, 39, 72], which suggest that even under the same pretrained feature space, functionally similar features can be manipulated into different arrangements throughout the course of finetuning.
@@ 4.1 Hypothesis
Building on the above observations, we know that weight delta entries at identical positions do not necessarily encode consistent semantics or functionalities. This invalidates direct entry-wise interference reduction of the weight delta matrices. Therefore, we make the following hypothesis: weight deltas of differently adapted models must be decomposed and then coordinated into an identical representation space, before proper interference reduced merging can be carried out.
@@ 4.2 Algorithm
We propose Decom-Renorm-Merge (DRM) which comprises (1) shared decomposition via SVD, for aligning different weight deltas into a joint representation space, (2) per-task renormalization for stabilizing each individual basis and enabling downstream interference reductions, (3) TIES’ operations applied in the Decompose-renormalized joint space, including pruning, sign election, and disjoint averaging. See Appendix E for analysis on different choice of decompositions.
@@
4) Pruning: For each task $t$, we keep only the top-$k\%$ largest-magnitude entries in $\tilde{V}^{T}_{t}$, and zero out the others. It is vital that this step be done after renormalization to ensure stability, pruning on unnormalized magnitudes could distort the basis vectors. We default this to top-$20\%$ [83].
5) Sign Election: To resolve inter-task sign conflicts, we compute a sign mask, $S=\operatorname{sgn}(\sum_{t=1}^{N}\Sigma_{t}\tilde{V}_{t}^{T})$, selecting a magnitude dominant sign for each position. Then, we zero out entries with signs opposite to the elected mask. This procedure is done in scale-aware fashion by multiplying the $\Sigma_{t}$ in.
@@
Here, $\lambda_{t}$ is a task-specific scalar weighting coefficient, with a default $\lambda=1.0$ for all tasks. $\Gamma$ is an entry-wise averaging reciprocal mask that ensures averaging procedure is performed only over non-zero entries. We define $\Gamma\in\mathbb{R}^{m\times n}$ with entry at row $i$, column $j$ as the following reciprocal:
@@ Methods
For merging of large language models, we adapt Llama3.1-8B [76] to five GLUE tasks individually: MNLI [80], QNLI [78], RTE [78], COLA [79], and SST2 [70] For this setup, we perform low-rank adaptation ($r=16$) instead of full finetuning as done previously. Table 2 shows that without validation data (no hyperparameter tuning), DRM-H outperforms other competitors by more than $1.9\%$. In the presence of hyperparameter tuning, DRM-V outperforms the strongest baseline by a margin of $3.9\%$. Notice that, unlike in the previous winning scenarios, where DRM-V is usually subpar or relatively close to DRM-H. Reversed here, DRM-V with hyperparameter tuning surpasses DRM-H by a margin of $1.0\%$. We provide an analysis on this phenomenon in Appendix G.
@@ 6 Understanding the Decompose-Renormalized Joint Space
Seeing the performance of the DRM, we hope to understand more about our method, particularly the intrinsic nature of the Decompose-renormalized joint space used for the merging.
We perform a series of studies to answer the following questions: 1) Why is it crucial that magnitude pruning be done on renormalized singular vector matrices? Shouldn’t pruning on unrenormalized basis vector matrices be even better? Since they better capture the real scale of each basis entry. 2) How does reducing interference on the Decompose-renormalized space reflect interference on the original parameter space? Does reducing interference in one space directly influence the other? 3) How large is the difference between individual task and the joint decomposed weight delta?
@@
6.1 Why is it Crucial that Weight Pruning be Done on the Renormalized Singular Vector Matrices?
After partitioning the right singular vector matrix $V^{T}$ of the concatenated SVD, the row basis vectors lose the orthonormal property. We renormalize each row back to unit norm [^2] Unlike the unit length property, orthogonality of the split matrices, is not easily reattainable. We observe that enforcing orthogonality through Gram-Schmidt process harshly degrades the performance. before applying interference reduction techniques. As shown in Section 5.3 that renormalization is central to DRM in terms of performance, here we investigate the effect of renormalization on the pruning behavior, from both empirical and theoretical perspectives.
Empirical investigation.
Figure 3 displays the percentages of entries being dropped from each basis vector. We see that without renormalization, the dropping densities fluctuate severely, with some vectors have almost all of their entries zeroed out. Removing top principle components like this would quickly deteriorate downstream performance. By adding in renormalization, the dropping densities are stabilized evenly across all vectors, bounded to around the configured prune rate of 50%, no single vector is wholly pruned.
Theoretical investigation.
The following proposition presents a theoretical explanation for the empirically observed pruning fluctuation when renormalization is not applied.
Proposition 1 (Shared Norm Budget of Partitioned Basis Vectors).
Let $V^{T}$ from a horizontally concatenated SVD be partitioned into task-specific blocks, e.g. $V^{T}=[V_{A}^{T}\;\;V_{B}^{T}]$ for two tasks. For any given row index $i$, the squared norms of the corresponding row vectors $v_{A,i}^{T}$ and $v_{B,i}^{T}$ must satisfy $\|v_{A,i}^{T}\|_{2}^{2}+\|v_{B,i}^{T}\|_{2}^{2}=1$. This creates a "unit budget" shared between tasks for each basis direction. Consequently, a larger norm for one task’s vector necessitates a smaller norm for the others.
The proof and formal statement are deferred to Appendix D.1. Informally, this unit budget biases magnitude-based pruning to favor preserving entries from the vector with the larger norm, as in the absence of outliers, its entries will have a larger magnitude on average. This can lead to disproportionate pruning of basis vectors associated with tasks that have a smaller share of the norm budget, regardless of the intrinsic importance of their features. Renormalization corrects this pruning bias by equalizing norms of the partitioned basis vectors.
Table 4: Average sign agreement after pruning different spaces. We compare between pruning in the original parameter space (labeled Original Weight), versus pruning in the joint singular space. For latter, sign agreement is measured in both the joint singular space itself (labeled DS), and after projecting back to the original space (labeled OS). We find that reducing interference on the joint space does not correspond to equally high sign agreement in the original parameter space. Note that the minimum sign agreement is $50\%$, when there are an equal number of positive and negative entries.
@@ 6.2 How Does Pruning on Decompose-renormalized Space Reflect Interference in Original Parameter Space?
However, our analysis in Table 4 reveals that although pruning in the Decompose-renormalized space effectively improves merging performance, it does not result in equally large sign agreement when projected back to the original parameter space. Specifically, we first project the weight deltas into the Decompose-renormalized space, apply pruning to the basis entries. We then transform the results back to the original space and measure sign agreement, but we found the agreement on the original space to be considerably less.
Nevertheless, DRM still achieves mostly superior merging performance compared to methods reducing interference directly in the original parameter space. This suggests that interference reduction techniques lead to the most effective merging when applied in the Decompose-renormalized space, rather than directly on individual parameter entries. See Appendix F for more detailed results and breakdown on different agreement levels.
6.3 How Large is the Difference between Individual Task and the Joint Decomposed Weight Delta?
We theoretically evaluate the difference between weight delta of each individual task, and the joint decomposed weight delta used in DRM-H. We express the difference through their corresponding singular values (i.e. $i$-th singular value of a single weight delta matrix, and $i$-th singular value of the concatenated matrix), and establish the following bound:
Proposition 2 (Bounded Difference of Weight Delta Concatenation).
Suppose that $t$ is an arbitrary task index, and $M\in\mathbb{R}^{m\times kn}$ is a concatenation of $k$ copies of $\Delta W_{t}$, where $k$ is the total number of tasks and $r$ is the rank of $M$. Let $\widetilde{M}=[\Delta W_{1}\;\dots\;\Delta W_{k}]$ be the concatenation of the collection $\{\Delta W_{j}\}_{j=1}^{k}$. We define the perturbation matrix
[…]
Then, the difference between $i$-th singular value of the concatenated matrix $\widetilde{M}$ and of the individual $\Delta W_{t}$, for all $i\in[1,r]$ and $t\in[1,k]$, is bounded by
$|\sigma_{i}(\widetilde{M})-\sqrt{k}\sigma_{i}(\Delta W_{t})|\leq\frac{1}{\sqrt{k}\sigma_{i}(\Delta W_{t})}\sum_{j=1}^{k}\left(2\|\Delta W_{t}\|_{2}\|E_{j}\|_{2}+\|E_{j}\|_{2}^{2}\right).$
This result quantifies the bounded behavior of change to each individual task’s weight delta, such that the alteration induced by horizontal concatenation of weight deltas will never exceed a specific amount described by the bound. It is evident that larger singular value $\sigma_{i}(\Delta W_{t})$ will exhibit a tighter bound due to the reciprocal term $\frac{1}{\sigma_{i}(\Delta W_{t})}$ on the right hand side. The proof and additional discussion are provided in Appendix D.2.
Intuitive interpretation.
When comparing jointly decomposed weight deltas of different models during interference reduction, it is essential to preserve the top principle features of each model to ensure a fair comparison across tasks; otherwise, certain tasks could be degraded unknowingly. Our derived bound addresses this by indicating a bounded degree of change, especially for the important feature bases characterized by high singular values. It assures the preservation of salient information from each model, facilitating interference reduction and merging in the subsequent steps.
@@ 8 Acknowledgements
We are grateful to Artit Suwanbandit and Patawee Prakrankamanant for inspiration and insightful discussions early into this work. Thanks also to Chompakorn Chaksangchaichot for helpful comments that improved the presentation of this paper.
@@ Appendix C Significance Analysis of the Merging Performance
Appendix D Theoretical Details
D.1 Biased Magnitude-based Pruning of Partitioned Singular Basis
Proposition 3 (Shared Norm Budget of Partitioned Orthonormal Basis Vectors).
Let $V\in\mathbb{R}^{kd\times r}$ be a matrix with orthonormal columns, such that $V^{T}V=I_{r}$. Let $V^{T}$ be partitioned horizontally into $k$ blocks corresponding to $k$ tasks, $V^{T}=[V_{1}^{T}\;\;V_{2}^{T}\;\dots\;V_{k}^{T}]$, where each block $V_{t}^{T}\in\mathbb{R}^{r\times d}$. Let $v_{t,i}^{T}$ denote the $i$-th row vector of block $V_{t}^{T}$.
Then, for any given row index $i\in\{1,\dots,r\}$, the sum of the squared Euclidean norms of the corresponding row vectors across all task blocks is equal to one:
$\sum_{t=1}^{k}\|v_{t,i}^{T}\|_{2}^{2}=1\,.$
This demonstrates that the squared norms of the partitioned row vectors for any basis direction $i$ are constrained by a shared unit budget.
Proof.
We begin with the orthonormality property of the matrix $V$:
$V^{T}V=I_{r}\,.$
Substituting the partitioned form of $V^{T}$ and its corresponding transpose $V$:
$V_{1}^{T}\;\;V_{2}^{T}\;\dots\;V_{k}^{T}]\begin{bmatrix}V_{1}\\ V_{2}\\ \vdots\\ V_{k}\end{bmatrix}=\sum_{t=1}^{k}V_{t}^{T}V_{t}=I_{r}\,.$
This equation establishes an identity between the sum of matrix products and the identity matrix $I_{r}$. We are interested in the diagonal entries of this identity. The $i$-th diagonal entry is given by:
$\left(\sum_{t=1}^{k}V_{t}^{T}V_{t}\right)_{ii}=\sum_{t=1}^{k}(V_{t}^{T}V_{t})_{ii}=(I_{r})_{ii}=1\,.$
The term $(V_{t}^{T}V_{t})_{ii}$ represents the inner product of the $i$-th row of $V_{t}^{T}$ (which is $v_{t,i}^{T}$) with the $i$-th column of $V_{t}$. Since the $i$-th column of $V_{t}$ is the transpose of the row vector $v_{t,i}^{T}$, this inner product is equivalent to the squared Euclidean norm of the row vector:
$(V_{t}^{T}V_{t})_{ii}=v_{t,i}^{T}(v_{t,i}^{T})^{T}=\|v_{t,i}^{T}\|_{2}^{2}\,.$
Substituting this back into the sum gives the final result:
$\sum_{t=1}^{k}\|v_{t,i}^{T}\|_{2}^{2}=1\,.$
This holds for all row indices $i=1,\dots,r$. ∎
Due to the existence of this constraint, a larger norm for one task’s vector necessitates a smaller norm for the others. The connection from this norm budget to the entry magnitude is given in the next corollary.
Corollary 3.1 (Larger RMS Magnitude in Vector with Larger Norm Share).
An unequal distribution of this norm share (i.e. $\|v_{A,i}^{T}\|_{2}^{2}\neq\|v_{B,i}^{T}\|_{2}^{2}$ for two tasks $A$ and $B$) directly implies that the root-mean-square (RMS) magnitude of entries is higher in the vector with the larger norm share.
Proof.
Consider two task-specific $d$-dimensional row vectors, $v_{A,i}^{T}$ and $v_{B,i}^{T}$, for a given basis direction $i$. Assume without loss of generality that task A has a larger norm share for this direction, such that
$\|v_{A,i}^{T}\|_{2}>\|v_{B,i}^{T}\|_{2}\;.$
Since the norms are non-negative, squaring both sides preserves the inequality:
$\|v_{A,i}^{T}\|_{2}^{2}>\|v_{B,i}^{T}\|_{2}^{2}\;.$
The root-mean-square (RMS) magnitude of a vector $v$ of dimension $d$ is defined as $\text{RMS}(v)=\sqrt{\frac{1}{d}\sum_{j=1}^{d}v_{j}^{2}}=\frac{\|v\|_{2}}{\sqrt{d}}$.
Applying this definition to our two vectors gives
$\text{RMS}(v_{A,i}^{T})=\frac{\|v_{A,i}^{T}\|_{2}}{\sqrt{d}}\quad\text{and}\quad\text{RMS}(v_{B,i}^{T})=\frac{\|v_{B,i}^{T}\|_{2}}{\sqrt{d}}\;.$
Given our initial assumption that $\|v_{A,i}^{T}\|_{2}>\|v_{B,i}^{T}\|_{2}$, it directly follows that:
$\text{RMS}(v_{A,i}^{T})>\text{RMS}(v_{B,i}^{T})\;.$
Thus, a larger norm share necessarily leads to a higher RMS magnitude. ∎
Under the reasonable assumption that the entries of $v_{A,i}^{T}$ and $v_{B,i}^{T}$ are drawn from distributions of the same shape, their scale will be determined by their respective RMS magnitudes. Consequently, the entries of the vector with the lower norm share are statistically smaller. Therefore, applying a uniform magnitude-based pruning across these un-renormalized vectors will disproportionately remove entries from the vector with the lower norm share.
D.2 Bound of Difference Between Individual Task and the Concatenated Matrix
This section provides derivation and formal discussion of the Proposition 2 (Bounded difference of weight delta concatenation) from Section 6.3. We begin by stating a theorem proven by Shamrai [68]: